# UEMM-Air: A Synthetic Multi-modal Dataset for Unmanned Aerial Vehicle Object Detection

### v1

@@ 4.1 Experimental Setup
Benchmark Models. We adopted YOLOv8-L [12], FasterRCNN [26] with ResNet50 [10], RT-DETR-L [41] as the general object detection baseline models. In the multi-modal object detection experiment, we designed a dual-path multi-modal detector with mid-level feature fusion. We utilized YOLOv7-L [36] as the base detector, with two separate backbone networks to extract features from two modalities. We also designed a feature fusion module that utilizes Coordinate Attention (CA) [11]. Specifically, we first directly concatenated the features of two modalities and employed Coordinate Attention to fuse them simultaneously in terms of channel and spatial information. The fused features were then entered into the neck part of the detector to complete the remaining detection tasks.
| Method | LargeVehicle | MediumVehicle | SmallVehicle | |
| YOLOv8 | 55.6% | 73.7% | 75.9% | 68.4% |
| FasterRCNN | 32.4% | 41.5% | 31.2% | 34.9% |
| RT-DETR | 49.8% | 55.8% | 77.5% | 61.1% |
Training Settings. All experiments were conducted in Pytorch with a NVIDIA RTX 3090 GPU. During the model transferability verification, we set the batch size to 16 and trained for 200 epochs. In other experiments, we froze the backbone network of the detector and trained for 50 epochs with a batch size of 32. All detectors were trained using an Adam optimizer [13] with a momentum of 0.937. The learning rate was initialized as 0.001 with a cosine decay [19]. We fix random seed to 18 to ensure the experiment’s reproducibility.
4.2 Transferability Verification
To demonstrate the advantage of our UEMM-Air on model transferability, We pre-trained two detectors utilizing SynDrone and UEMM-Air, respectively. Then we subsequently fine-tuned the models on the VisDrone and UAVDT datasets. The experimental results are presented in Table 2. While the number of images is smaller than SynDrone (20 & 60), the model pre-trained on the UEMM-Air dataset demonstrates stronger generalization performance on real-world scenario data. For example, after obtaining pre-trained weights on UEMM-Air and SynDrone datasets, we fine-tuned the YOLOv8 model on the VisDrone dataset. The model pre-trained on UEMM-Air demonstrated a 2.7% improvement in and a 1.7% improvement in . This might be attributed to the provision of more accurate annotations, more categories, and more diverse scenarios in UEMM-Air for the model pre-training process.
4.3 Evaluation on Object Detection
| Method | BoxTruck | Bulldozer | Excavator | ForkLift | Jeep | Motor | Pickup |
| YOLOv8 | 61.3% | 62.5% | 42.2% | 51.0% | 72.6% | 75.9% | 78.7% |
| FasterRCNN | 33.9% | 51.5% | 19.4% | 24.4% | 40.6% | 30.7% | 50.6% |
| RT-DETR | 69.8% | 58.2% | 55.9% | 61.2% | 23.1% | 79.8% | 81.8% |
| Method | SUV | Trailer | Truck | Van | RoadRoller | Sedan | |
| YOLOv8 | 81.9% | 51.1% | 56.1% | 69.0% | 56.7% | 71.6% | 63.9% |
| FasterRCNN | 46.1% | 30.6% | 29.5% | 35.9% | 39.6% | 40.1% | 36.4% |
| RT-DETR | 85.1% | 44.3% | 58.5% | 79.2% | 63.2% | 79.4% | 64.6% |
General Object Detection. We employed coarse-grained and fine-grained labels for training three detectors, namely YOLOv8, FasterRCNN, and RT-DETR, respectively. Experimental outcomes are presented in Table 3 and Table 4. In the coarse-grained detection experiments, YOLOv8 achieved the highest precision, with an of 68.4%, followed by RT-DETR with a precision of 61.1%. In the fine-grained detection experiments, RT-DETR exhibited the highest at 64.6%. Across both experiments, FasterRCNN consistently demonstrated lower performance compared to the other two detectors, consistent with the performance of the three detectors in the general domain.
| Modality | ||
| RGB | 68.4% | 53.2% |
| RGB + Seg | 75.3% | 54.1% |
| RGB + Surface | 74.3% | 54.4% |
| RGB + Depth | 73.0% | 54.2% |
Multi-modal Object Detection. In Table 5, we conducted mid-level fusion experiments for multi-modal object detection with RGB modality and other three modalities. The model fusion of RGB with segmentation modality achieved the best performance on , surpassing the baseline model (RGB only) by 6.9%. The fusion of RGB with surface normal modality achieved the best performance on , surpassing the baseline model by 1.2%. However, fusion with depth modalities resulted in the lowest performance. This could be due to the distinct features of object positions in segmentation modality and the detailed texture features in surface normal, both containing more effective information compared to depth.
@@ 4.3 Evaluation on Object Detection
4.4 Evaluation on Automatic Annotation Algorithm
@@
Verification of the Annotations’ Reliability. We randomly divided the dataset into 5 parts and conducted five-fold cross-validation experiments on the YOLOv8 model. The experimental results are demonstrated in Table 6. By sequentially using different sets of 4 parts as the training set and the remaining part as the validation set, we observed that the results of the 5-fold cross-validation were quite similar. The lowest was 62.3%, and the highest was 64.3%, with a range of 2.0%. This result indicates that the annotations we generated are consistent in their distribution.
@@
Comparison with Manual Annotations. We randomly selected a subset of images for manual annotation and then visually compared them with our automatic annotations. As presented in Fig 5, we visualized our generated labels and manually annotated labels separately for comparison. It can be observed that our annotations are almost identical in position to the manual labels. Moreover, our annotation algorithm has some advantages in labeling small objects. We found that when objects are far away, manual annotations may contain errors due to the smaller scale of the objects. For example, manual annotations may not be as closely aligned with the edges of the objects as our generated annotations. It will introduce more foreground information, which could impact the model’s accuracy.
5 Conclusion
@@
In this paper, we released a synthetic UAV-based object detection dataset, named UEMM-Air. Our work achieved three main breakthroughs. Firstly, to the best of our knowledge, UEMM-Air is the UAV-OD dataset with the largest number of aligned modalities (RGB, segmentation, depth, surface normals, and IMU parameters). Secondly, we designed a new automatic annotation method to for UAV-OD tasks, enhancing the accuracy of annotations by employing segmentation and depth images. Finally, we conducted a series of experiments. The results validated that the models pre-trained on UEMM-Air exhibit strong transferability. We also established new benchmarks across various types of UAV-OD tasks, including fine-grained object detection and multi-modal object detection. We will continue to build new simulated scenarios in the future to expand the scale and number of modalities in our dataset, supporting research on UAV vision tasks.
Broader Impacts and Limitations
Broader impacts. Our UEMM-Air dataset features 5 aligned modalities, making it the dataset with the highest number of aligned modalities known to date. This will effectively support UAV multi-modal object detection tasks. Furthermore, leveraging the advantage of easily accessible generative data, the paired multi-modal data we provide will also be beneficial for supporting large-scale multi-modal pre-training tasks.
Limitations. Due to the simulated scenes still requiring manual construction, there are still certain manual costs associated with building the dataset. However, the development of scene dialogue generation tasks in the Unreal Engine is already under research. We believe that it won’t be long before automated scene construction is achieved, integrating with our construction methods to efficiently generate vast amounts of data for model pre-training.
@@
This work was partially supported by the Fundamental Research Funds for the Central Universities (No. B240201077), National Nature Science Foundation of China (No. 62372155 and No. 62302149), Aeronautical Science Fund (No. 2022Z071108001), Joint Fund of Ministry of Education for Equipment Pre-research (No. 8091B022123), Water Science and Technology Project of Jiangsu Province under grant No. 2021063, Qinglan Project of Jiangsu Province, Postgraduate Research & Practice Innovation Program of Jiangsu Province (No. SJCX24_0183).

### v2

@@ IV-A1 Benchmark Models
We adopted YOLOv11 [67], RT-DETR [68] and D-FINE [69] as the general object detection baseline models. In the multi-modal object detection experiments, we designed a dual-path multi-modal detector with mid-level feature fusion. We utilized YOLOv7-L [70] as the base detector, with two separate backbone networks to extract features from two modalities. We also designed a feature fusion module that utilizes Coordinate Attention (CA) [71]. Specifically, we first directly concatenated the features of two modalities and employed Coordinate Attention to fuse them simultaneously in terms of channel and spatial information. The fused features were then entered into the neck part of the detector to complete the remaining detection tasks.
@@
In image-text contrastive learning experiments, we employed OpenAI CLIP [27] and RemoteCLIP [66] as benchmark models. We selected four types of visual backbone architecture for the two CLIP models, ranging from ResNet-50, ViT-Base-16, ViT-Base-32, and ViT-Large-14. We utilized the transformer architecture, consisting of 12 layers and 8 attention heads for text encoder. The maximum token sequence length is set to 77. The InfoNCE [72] loss operates on the [CLS] token produced by the image and text backbone.
In referring image segmentation experiments, we adopted three SoTA transformer architecture models: LAVT [73], RMSIN [74], RefSegformer [75]. Among them, RMSIN is specially designed for remote sensing scenarios.
| Method | Fine-tuned | Pre-trained | ||
| YOLOv11 [67] | VisDrone | SynDrone | 25.6% | 16.0% |
| UEMM-Air | 28.3% | 17.7% | ||
| UAV-DT | SynDrone | 85.4% | 55.6% | |
| UEMM-Air | 86.2% | 56.1% | ||
| FasterRCNN [44] | VisDrone | SynDrone | 5.1% | 1.1% |
| UEMM-Air | 5.5% | 2.3% | ||
| UAV-DT | SynDrone | 48.0% | 9.5% | |
| UEMM-Air | 53.8% | 14.7% |
IV-A2 Training Settings
All detection and segmentation experiments were conducted in Pytorch with a NVIDIA RTX 3090 GPU. During the model transferability verification, we set the batch size to 16 and trained for 200 epochs. In other object detection experiments, we froze the backbone network of the detector and trained for 50 epochs with a batch size of 32. All detectors were trained using an Adam optimizer [76] with a momentum of 0.937. The learning rate was initialized as 0.001 with a cosine decay [77]. We fix random seed to 18 to ensure the experiment’s reproducibility.
The image-text contrastive learning experiments were trained on an NVIDIA RTX 3090 GPU. The training process was accelerated by employing the Adam optimizer [76]. The learning rate was set to 7e-5, 4e-5, and 1e-4, respectively, for ResNet-50, ViT-Base-32, and ViT-Large-14 models, and the corresponding batch size was set to 256, 128, and 28, respectively.
The referring image Segmentation experiments were deployed in 4 NVIDIA RTX 4090 GPUs. The initial learning rate was set to 0.0003, with a batch size of 6 per GPU and the Adam [76] optimizer. The training was conducted for a total of 40 epochs. The input size of images was set to 480 480.
IV-B Evaluation on Automatic Annotation Algorithm
@@
| Method | Version | Input Size | #Params (M). | BFLOPs | FPS (bs=1) | ||||||
| YOLOv11 [67] | n | 640×640 | 50.8% | 68.9% | 54.9% | 22.0% | 80.6% | 94.5% | 2.6 | 6.3 | 263.2 |
| s | 640×640 | 56.3% | 73.8% | 61.5% | 29.2% | 84.4% | 96.3% | 9.4 | 21.3 | 212.8 | |
| m | 640×640 | 57.0% | 74.8% | 62.3% | 30.4% | 84.6% | 96.0% | 20.0 | 67.7 | 185.2 | |
| l | 640×640 | 58.6% | 76.0% | 64.6% | 32.8% | 85.5% | 96.8% | 25.3 | 86.6 | 158.7 | |
| x | 640×640 | 60.0% | 76.8% | 66.1% | 34.7% | 86.5% | 97.3% | 56.8 | 194.5 | 94.3 | |
| RT-DETR [68] | l | 640×640 | 38.4% | 60.3% | 40.3% | 11.0% | 64.1% | 84.7% | 32.0 | 103.5 | 90.9 |
| x | 640×640 | 40.1% | 73.7% | 38.4% | 15.7% | 63.1% | 82.4% | 65.5 | 222.5 | 56.2 | |
| D-FINE [69] | n | 640×640 | 63.5% | 82.5% | 68.7% | 41.9% | 86.3% | 95.2% | 7.1 | 3.7 | 44.1 |
| s | 640×640 | 71.8% | 90.3% | 78.0% | 54.0% | 89.8% | 96.2% | 24.9 | 10.2 | 38.8 | |
| m | 640×640 | 73.2% | 90.7% | 78.8% | 55.6% | 91.4% | 97.1% | 56.4 | 19.2 | 31.4 | |
| l | 640×640 | 74.3% | 91.3% | 80.0% | 57.2% | 91.8% | 97.3% | 90.7 | 30.7 | 23.1 | |
| x | 640×640 | 75.8% | 92.4% | 81.8% | 59.5% | 92.5% | 97.5% | 61.55 | 202.2 | 18.8 |
IV-B1 Verification of the Annotations’ Reliability
@@
We randomly divided the dataset into 5 parts and conducted five-fold cross-validation experiments on the YOLOv11 model. The experimental results are demonstrated in Table II. By sequentially using different sets of 4 parts as the training set and the remaining part as the validation set, we observed that the results of the 5-fold cross-validation were quite similar. The lowest was 62.3%, and the highest was 64.3%, with a range of 2.0%. This result indicates that the annotations we generated are consistent in their distribution.
| Method | Version | Input_Size | |||
| YOLOv11 | n | 640×640 | 31.4% | 58.8% | 30.2% |
| s | 640×640 | 34.8% | 64.3% | 33.9% | |
| m | 640×640 | 37.7% | 67.9% | 37.1% | |
| l | 640×640 | 38.3% | 68.6% | 37.9% | |
| x | 640×640 | 39.1% | 69.5% | 38.9% |
| Modality | |||
| RGB | 50.3% | 68.4% | 53.2% |
| RGB + Seg | 55.6% | 75.3% | 54.1% |
| RGB + Surface | 55.8% | 74.3% | 54.4% |
| RGB + Depth | 53.5% | 73.0% | 54.2% |
| RGB + Seg + Surface + Depth | 57.3% | 78.2% | 58.0% |
IV-B2 Comparison with Manual Annotations
@@
We randomly selected a subset of images for manual annotation and then visually compared them with our automatic annotations. As presented in Fig. 10, we visualized our generated labels and manually annotated labels separately for comparison. It can be observed that our annotations are almost identical in position to the manual labels. Moreover, our annotation algorithm has some advantages in labeling small objects. We found that when objects are far away, manual annotations may contain errors due to the smaller scale of the objects. For example, manual annotations may not be as closely aligned with the edges of the objects as our generated annotations. It will introduce more foreground information, which could impact the model’s accuracy.
IV-C Evaluation on Object Detection
IV-C1 Transferability Verification
To demonstrate the advantage of our UEMM-Air on model transferability, We pre-trained two detectors utilizing SynDrone and UEMM-Air, respectively. Then we subsequently fine-tuned the models on the VisDrone and UAVDT datasets. The experimental results are presented in Table III. While the number of images is smaller than SynDrone (20 & 60), the model pre-trained on the UEMM-Air dataset demonstrates stronger generalization performance on real-world scenario data. For example, after obtaining pre-trained weights on UEMM-Air and SynDrone datasets, we fine-tuned the YOLOv8 model on the VisDrone dataset. The model pre-trained on UEMM-Air demonstrated a 2.7% improvement in and a 1.7% improvement in . This might be attributed to the provision of more accurate annotations, more categories, and more diverse scenarios in UEMM-Air for the model pre-training process.
IV-C2 General Object Detection
We trained several state-of-the-art detection frameworks, including YOLOv11 [67], RT-DETR [47], and D-FINE [69], utilizing our UEMM-Air. Experimental outcomes are presented in Table IV. In terms of detection accuracy, the D-FINE-x model achieved the best performance with a mean Average Precision (mAP) of 75.8%. Similar to performances on other UAV-OD datasets, the detection accuracy for small objects was 59.5%, which is significantly lower compared to 97.5% for large objects. This indicates that our dataset poses significant challenges for small object detection, providing valuable insights for researchers aiming to tackle the difficulties associated with small object detection. Additionally, considering the real-time requirements of UAVs, we also tested the inference time metrics. The YOLOv11-n, as a lightweight model, achieved 263.2 FPS on an RTX 3090 GPU, with only 2.6M parameters.
| Method | publication | Image Backbone | Text Backbone | Image To Text | Text To Image | Mean Recall | ||||||
| R@1 | R@5 | R@10 | MR | R@1 | R@5 | R@10 | MR | |||||
| CLIP [27] | ICML2021 | ResNet50 | Transformer | 6.79 | 27.38 | 44.05 | 26.07 | 4.57 | 19.33 | 32.41 | 18.77 | 22.42 |
| ViT-B-16 | 8.67 | 30.77 | 47.35 | 28.93 | 5.96 | 22.48 | 35.71 | 21.38 | 25.15 | |||
| ViT-B-32 | 12.80 | 43.51 | 62.72 | 39.59 | 8.59 | 31.09 | 47.82 | 29.17 | 34.38 | |||
| ViT-L-14 | 12.81 | 44.62 | 61.99 | 39.32 | 8.68 | 30.68 | 46.69 | 29.08 | 34.16 | |||
| RemoteCLIP [66] | TGRS2024 | ResNet50 | Transformer | 7.02 | 27.34 | 43.72 | 26.03 | 4.56 | 19.17 | 32.33 | 18.69 | 22.36 |
| ViT-B-32 | 12.24 | 41.58 | 60.77 | 38.19 | 8.39 | 30.84 | 47.52 | 28.92 | 33.56 | |||
| ViT-L-14 | 12.31 | 41.21 | 60.84 | 38.12 | 8.36 | 30.91 | 46.63 | 28.46 | 33.29 | |||
| Methods | Publication | Image Backbone | Text Backbone | Pr@0.5 | Pr@0.6 | Pr@0.7 | Pr@0.8 | Pr@0.9 | oIoU | mIoU |
| LAVT [73] | CVPR22 | Swin-B | BERT | 53.17% | 42.42% | 35.47% | 23.69% | 8.03% | 63.72% | 51.09% |
| RMSIN [74] | CVPR24 | Swin-B | BERT | 57.40% | 49.20% | 38.80% | 25.40% | 8.60% | 65.82% | 51.97% |
| RefSegformer [75] | IEEE TIP24 | Swin-B | BERT | 55.28% | 48.55% | 38.01% | 26.04% | 8.62% | 64.99% | 51.23% |
IV-C3 Multi-modal Object Detection
In Table VI, we conducted mid-level fusion experiments for multi-modal object detection with RGB modality and the other three modalities. The model fusion of RGB with segmentation modality achieved the best performance on , surpassing the baseline model (RGB only) by 6.9%. The fusion of RGB with surface normal modality achieved the best performance on , surpassing the baseline model by 1.2%. However, fusion with depth modalities resulted in the lowest performance. This could be due to the distinct features of object positions in segmentation modality and the detailed texture features in surface normal, both containing more effective information compared to depth. We also combined 4 modalities in our experiments, achieving a mAP of 57.3%, which is an improvement of 7% compared to utilizing the RGB modality alone.
IV-D Evaluation on Instance Segmentation
We selected the YOLOv11 framework to conduct instance segmentation experiments, as shown in Table V. It provides benchmark results for our dataset. As the largest scale model, YOLOv11-x achieved the best accuracy, with a of 39.1%, of 69.5%, and of 38.9%. Compared to object detection, instance segmentation is a more challenging dense prediction task, resulting in a relatively lower .
IV-E Evaluation on Image-Text Contrastive Learning
Table VII presents the performance of CLIP and RemoteCLIP in image-text retrieval on our dataset. We report the retrieval recall of top-1 (R@1), top-5 (R@5), top-10 (R@10), and the mean recall of these values. From Table VII, it can be observed that the original OpenAI CLIP performs better, achieving the highest Mean Recall of 34.38. The versions using ViT-B-32 and ViT-L-14 as visual backbones perform similarly across several metrics, but ViT-B-32 demonstrates superior average performance in both retrieval tasks. Additionally, it is noteworthy that RemoteCLIP, as the CLIP model for remote sensing, performs worse than the original CLIP. This may be due to the reduced generalization capability of RemoteCLIP after fine-tuning on remote sensing satellite data.
IV-F Evaluation on Referring Segmentation
We conducted experiments with three state-of-the-art referring image segmentation models on our UEMM-Air, with the results summarized in Table VIII. The results indicate that RMSIN exhibits the best performance on both the mIoU and oIoU average performance metrics. It can be attributed to RMSIN’s architecture is specifically designed for aerial perspectives, making it more conducive to feature learning from a UAV’s viewpoint. Additionally, RMSIN’s metrics at Pr@0.8 and Pr@0.9 are slightly lower than those of RefSegformer, aligning with the performance differences reported in the original paper for the two models. The phenomenon indirectly validates the high accuracy of the annotations generated automatically in our dataset.
V Conclusion
@@
In this paper, we release a synthetic UAV-based environmental perception dataset, named UEMM-Air. Our work achieves three main breakthroughs: Firstly, to the best of our knowledge, UEMM-Air is the largest in terms of data scale, featuring the most paired modalities and the highest number of task types. Secondly, we design a new automatic annotation method, enhancing the accuracy of annotations by employing segmentation and depth images. Then, we generate a large number of text descriptions utilizing the annotations, further enriching our dataset with text modality. Finally, we provide benchmark results across multiple tasks, thereby expanding the breadth of tasks in the field of UAV-based environmental perception. We will continue to build new simulated scenarios in the future to expand the scale and number of modalities in our dataset, supporting research on UAV-based multi-modal perception tasks.
@@
This work was supported in part by the National Natural Science Foundation of China under Grant 62372155 and Grant 62302149, in part by the Postgraduate Research and Practice Innovation Program of Jiangsu Province under Grant SJCX24_0183, in part by the Fundamental Research Funds for the Central Universities under Grant B240201077, in part by the Aeronautical Science Fund under Grant 2022Z071108001, in part by the Qinglan Project of Jiangsu Province, and in part by Changzhou Science and Technology Bureau Project No. 20231313.